# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ Appendix B Regularity assumptions and expansivity results
To analyze the stability of iterative optimization algorithms, we typically rely on regularity assumptions on the loss function. We begin by stating standard regularity assumptions [23] used throughout the paper, along with classical results that follow from them.
@@
Notably, convex and $L$-smooth functions satisfy an important property known as co-coercivity (see Proposition 5.4 in [7]). Specifically,
@@
Here, for each $i\in\mathcal{H}$, $g^{(i)}$ denotes the exact gradient $\nabla\ell(\theta)$ in the case of $\mathrm{GD}$, or an unbiased estimate thereof—effectively computed with a sample from the local dataset—in the case of $\mathrm{SGD}$. We recall the expansiveness definition for an update rule from [23, Definition 2.3]:
@@
A straightforward extension of [23, Lemma 3.6] to the distributed setting yields:
@@
Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ and $\gamma>0$. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z):\Theta\to\mathbb{R}$ is L-smooth, Then $G^{\mathcal{A}}_{\gamma}$ is $\eta_{G^{\mathcal{A}}_{\gamma}}$-expansive [23, Definition 2.3], with the following expressions:
@@
Appendix C Robust aggregation rules
We present the Component-Wise Trimmed Mean $\mathrm{(CWTM)}$ and the Smallest Maximum Eigenvalue Averaging $\mathrm{(SMEA)}$ that we use throughout the paper. They are, respectively, instances of $(f,\kappa,\operatorname{Tr})$-robust and $(f,\kappa,\|\cdot\|_{sp})$-robust aggregation rules, each with an asymptotically optimal robustness parameter. Note that $(f,\kappa,\operatorname{Tr})$-robustness is computationally cheaper to satisfy and yields optimal convergence error in settings where the honest gradients’ covariance matrix has uniformly bounded trace over the parameter space. On the other hand, $(f,\kappa,\|\cdot\|_{sp})$-robustness is computationally expensive to satisfy—as robust aggregation rules must effectively operate in high-dimensional spaces, unlike simpler methods such as $\mathrm{CWTM}$—but yields optimal convergence error when only the spectral norm of the honest gradients’ covariance matrix is assumed to be bounded.
Definition C.1 (Smallest Maximum Eigenvalue Averaging $\mathrm{(SMEA)}$).
Given $f,n\in\mathbb{N}$, $f<n/2$ and vectors $g_{1},\ldots,g_{n}\in\mathbb{R}^{d}$, the smallest maximum eigenvalue averaging aggregation rule outputs:
$\overline{g}_{S^{*}}=\frac{1}{|S^{*}|}\sum_{i\in S^{*}_{t}}g_{i}$
with $\text{with }S^{*}\in\underset{S\subseteq\{1,\ldots,n\},|S|=n-f}{\text{argmin}}\lambda_{\max}\left(\frac{1}{|S|}\sum_{i\in S}\left(g_{i}-\overline{g}_{S}\right){\left(g_{i}-\overline{g}_{S}\right)}^{\intercal}\right)$
The $\mathrm{SMEA}$ aggregation rule is especially valuable because it is $(f,\kappa,\|\cdot\|_{sp})$-robust, with an optimal asymptotic robustness coefficient.
Lemma C.1 (Robustness of $\mathrm{SMEA}$ [5, Prop. 5.1]).
$\mathrm{SMEA}$ is $(f,\kappa,\|\cdot\|_{sp})$-robust with
$\kappa=\frac{4f}{n-f}{\left(1+\frac{f}{n-2f}\right)}^{2}.$
Assume that there exist $\nu$ constant such that $n\geq(2+\eta)f$, that is $f/n$ is strongly away from $1/2$, then $\kappa\in\mathcal{O}(f/n)$ is asymptotically optimal.
Unlike $\mathrm{SMEA}$, which leverages spectral properties to achieve robust aggregation, $\mathrm{CWTM}$ relies on coordinate-wise trimming.
Definition C.2 (Component-Wise Trimmed Mean $\mathrm{(CWTM)}$).
Given $f,n\in\mathbb{N}$, $f<n/2$, and real values $x_{1},\ldots,x_{n}$, we denote by $\sigma$ a permutation on $\{1,\ldots,n\}$ that sorts the values in non-decreasing order $x_{\sigma(1)}\leq\cdots\leq x_{\sigma(n)}$. Let $S_{x}=\{i\in[n];f+1\leq\sigma(i)\leq n-f\}$, the trimmed mean is given by:
$TM(x_{1},\ldots,x_{n})=\frac{1}{n-2f}\sum_{i=f+1}^{n-f}x_{\sigma(i)}=\frac{1}{n-2f}\sum_{i\in S_{x}}x_{i}$
We define the component-wise trimmed mean for vectors $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ as the coordinate-wise application of the trimmed mean operation. That is $CWTM(x_{1},\ldots,x_{n})$ is the vector whose $k-th$ coordinate is defined as
${[CWTM(x_{1},\ldots,x_{n})]}_{k}=TM({[x_{1}]}_{k},\ldots,{[x_{n}]}_{k})$
The $\mathrm{CWTM}$ aggregation rule is particularly useful as it is $(f,\kappa,\operatorname{Tr})$-robust, with an optimal asymptotic robustness coefficient.
Lemma C.2 (Robustness of $\mathrm{CWTM}$ [4, Prop. 2.1]).
$\mathrm{CWTM}$ is $(f,\kappa,\operatorname{Tr})$-robust with
$\kappa=\frac{6f}{n-2f}\left(1+\frac{f}{n-2f}\right).$
Assume that there exist $\nu$ constant such that $n\geq(2+\eta)f$, that is $f/n$ is strongly away from $1/2$, then $\kappa\in\mathcal{O}(f/n)$ is asymptotically optimal.
Appendix D Defered proofs from Section 3
This appendix contains the detailed proofs and technical results supporting the analysis presented in Section 3. We begin with supporting results in Section D.1, specifically, we bound the expectation of both the trace and the spectral norm of the empirical covariance matrix of honest workers’ gradients. Building on this, we establish bounds on uniform stability under Byzantine attacks in Section D.2, with detailed proofs of Theorems 3.1 and 3.4. These bounds are established for robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$ using any $(f,\kappa,\|\cdot\|_{op})$-robust aggregation rule, across strongly convex, convex, and nonconvex optimization settings. In Section D.3, we detail the properties of the $\mathrm{CWTM}$ aggregation rule, which is used to improve stability bound under poisoning attacks for smooth nonconvex objectives. Section D.4 focuses on leveraging loss function regularities under data poisoning, providing stability bounds for robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$ using the $\mathrm{SMEA}$ and $\mathrm{CWTM}$ rules across strongly convex, convex, and nonconvex settings. This corresponds to proofs of Theorems 3.2 and 3.5  Sections D.5 and D.6 details the unavoidable instability introduced by the $\mathrm{SMEA}$ aggregation rule by providing lower bounds on uniform stability under poisoning attacks for robust distributed $\mathrm{GD}$ and $\mathrm{SGD}$ across strongly convex and convex settings, that is proof of Theorem 3.3.
D.1 Expectation bound for the honest gradients’ empirical covariance trace and spectral norm
@@
We derive bounds on the expected trace and spectral norm of the empirical covariance matrix of the honest workers’ gradients under two settings the bounded gradient assumption (i.e. Lipschitz-continuity of the loss function).
Lemma D.1 (Expectation bound for the empirical covariance trace).
@@
Consider the setting described in Section 2 under Byzantine attacks. Assume the loss function $C$-Lipschitz. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa,\operatorname{Tr})$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
$\mathbb{E}_{\mathcal{A}}\operatorname{Tr}\Sigma_{\mathcal{H},t}=\mathbb{E}_{\mathcal{A}}[\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|g^{(i)}_{t}-\overline{g}_{t}\|_{2}^{2}]\leq C^{2}.$
Proof.
$\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|g^{(i)}_{t}-\overline{g}_{t}\|_{2}^{2}\leq\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}\|g^{(i)}_{t}\|^{2}_{2}\leq C^{2}.$
Lemma D.2 (Expectation bound for the empirical covariance spectral norm).
Consider the setting described in Section 2 under Byzantine attacks. Assume the loss function $C$-Lipschitz. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa,\|\cdot\|_{sp})$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
@@
D.2 Proofs of Theorems 3.1 and 3.4
In this section, we present proofs for Theorems 3.1 and 3.4. Specifically, we derive upper bounds on the uniform stability of learning algorithms under Byzantine attacks, considering convex, strongly convex, and nonconvex objectives. We begin with a unified proof sketch that outlines the key ideas common to both theorems. The proof of Theorem 3.1 is divided into two parts: one for $\mathcal{A}=\mathrm{GD}$, provided in Theorem D.3, and one for $\mathcal{A}=\mathrm{SGD}$, detailed in points (i) and (ii) of Theorem D.5. While the results are presented in a unified manner in the main text—due to their similar structure and identical bounds—we separate the proofs into two distinct components for clarity. The proof of Theorem 3.4 is established in point (iii) of Theorem D.5.
@@
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ with $F$ a $(f,\kappa,\|\cdot\|_{op})$-robust aggregation rule, and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. Let denote $m^{(i)}_{t}$ and $m^{\prime(i)}_{t}\in\mathbb{R}^{d}$ the local updates provided by worker $i\in\{1,\ldots,n\}$ along the trajectories $t\in\{0,\ldots,T-1\}$. In the context of the uniform algorithmic stability analysis via parameter sensitivity, we track how the parameters diverge along these optimization trajectories. To do so, we decompose the analysis by introducing an intermediate comparison between the robust update $G^{F}_{\gamma_{t}}$ and the averaging over honest workers update $G^{\mathcal{A}}_{\gamma_{t}}$. This enables us to leverage either the regularity of the loss function or the robustness property of the aggregation rule. This approach is primarily motivated by the fact that Byzantine vectors cannot be assumed to exhibit regularities when comparing them.
$\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]$ $=\mathbb{E}_{\mathcal{A}}[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})+G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})$
$\quad\quad\quad\quad\quad+G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]$
$\leq\underbrace{\mathbb{E}_{\mathcal{A}}\|G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}}_{A}$
$\quad+\underbrace{\gamma_{t}\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})\|_{2}+\|G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]}_{B}$
We then bound the term $A$ with classical stability tools from [23], and the term $B$ with the $(f,\kappa,\|\cdot\|_{op})$-robust property, the Jensen inequality and the Lemma D.1 (if $\|\cdot\|_{op}=\operatorname{Tr}$) or Lemma D.2 (if $\|\cdot\|_{op}=\|\cdot\|_{sp}$):
@@
D.2.1 Proof of Theorem 3.1 for GD
Theorem D.3 (Stability under Byzantine failure attacks - $\mathrm{GD}$).
Consider the setting described in Section 2 under Byzantine attacks. Let $\mathcal{A}=\mathrm{GD}$, with $F$ a $(f,\kappa,\|\cdot\|_{op})$-robust aggregation rule, with $\|\cdot\|_{op}\in\{\operatorname{Tr},\|\cdot\|_{sp}\}$, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
@@
For each $t\in\{0,\ldots,T-1\}$, following the proof sketch in D.2, we bound the term $\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ where $G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})$. For $t\in\{0,\ldots,T-1\}$, we have,
$\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ $=\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$
$\qquad+\frac{\gamma_{t}}{|\mathcal{H}|m}\|\nabla\ell(\theta_{t}^{\prime};z^{\prime(a,b)})-\nabla\ell(\theta_{t}^{\prime};z^{(a,b)})\|_{2}$
@@
where $\eta^{\mathrm{GD}}_{G{\gamma_{t}}}$ denotes the expansive coefficient, which depends on the regularity properties of the loss function (see Lemma B.2). Let define $\sigma^{F}_{\mathrm{GD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion by incorporating the impact of the Byzantine failure attack as demonstrated in the previous proof sketch.
@@
    We conclude the derivation of the bound by invoking the Lipschitz continuity assumption.∎
D.2.2 Proof of Theorem 3.1 for SGD and Theorem 3.4
Unlike $\mathrm{GD}$—where the differing samples affect every iteration—in $\mathrm{SGD}$, it may not be sampled immediately. This distinction enables our analysis to yield better stability upper bound with $\mathrm{SGD}$. Below, we state a lemma reasoning about the first occurrence of the differing samples.
Lemma D.4.
Consider the setting described in Section 2, under either Byzantine failure or poisoning attacks. We assume for all $z\in\mathcal{Z}$, $\ell(\cdot;z)$ nonnegative and uniformly bounded by $\ell_{\infty}\in\mathbb{R}_{+}$. Let $\mathcal{S}$ and $\mathcal{S}^{\prime}$ being two neighboring dataset, we denote $\theta_{T}=\mathcal{A}(\mathcal{S})$, $\theta_{T}^{\prime}=\mathcal{A}(\mathcal{S}^{\prime})$ and intermediate state of the algorithm $\theta_{t}$, $\theta_{t}^{\prime}$ and $\delta_{t}=\|\theta_{t}-\theta_{t}^{\prime}\|_{2}$ for any $t\in\{0,\ldots,T\}$. Then for every $z\in\mathcal{Z}$ and every $t_{0}\in\{0,\ldots,\min(m,T)\}$, the following holds for $\mathrm{SGD}$ with sampling with replacement:
$\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq\min\{\ell_{\infty},\frac{t_{0}}{m}\ell_{\infty}+\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)||\delta_{t_{0}}=0]\}$
Proof.
Let $z\in\mathcal{Z}$ and $t_{0}\in\{0,\ldots,\min(m,T)\}$, we have:
$\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]=\quad$ $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)||\delta_{t_{0}}=0]\mathbb{P}(\delta_{t_{0}}=0)$
$+\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)||\delta_{t_{0}}\neq 0]\mathbb{P}(\delta_{t_{0}}\neq 0)$
[…]
To bound $\mathbb{P}(\delta_{t_{0}}\neq 0)$, we consider $T_{0}$, the random variable of the first step the differing samples are drawn. We have:
$\mathbb{P}(\delta_{t_{0}}\neq 0)\leq\mathbb{P}(T_{0}\leq t_{0})\leq\min(1,\frac{t_{0}}{m})$
For the first inequality, we used the fact that under identical conditions (i.e., the same seed and initial state), the algorithm operating on two neighboring datasets will produce the same intermediate states until the differing samples are drawn. For the second inequality, we use the union bound to obtain $\mathbb{P}(T_{0}\leq t_{0})\leq\sum_{t=1}^{t_{0}}\mathbb{P}(T_{0}=t)=\min(1,\frac{t_{0}}{m})$. ∎
Building on the above lemma, we now state bounds for $\mathrm{SGD}$.
[…]
Consider the setting described in Section 2 under Byzantine attacks. Let $\mathcal{A}=\mathrm{SGD}$, with $F$ a $(f,\kappa,\|\cdot\|_{op})$-robust aggregation rule, with $\|\cdot\|_{op}\in\{\operatorname{Tr},\|\cdot\|_{sp}\}$, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
  • (i)
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S)};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq 2\gamma C^{2}T\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right).$
  • (ii)
  • (iii)
    If we assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ nonconvex, bounded by $\ell_{\infty}=\sup_{\theta\in\Theta,z\in\mathcal{Z}}\ell(\theta;z)$, with a monotonically non-increasing learning rate $\gamma_{t}\leq\frac{c}{Lt}$, $t\in\{0,\ldots,T-1\}$, $c>0$, we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
    $\sup_{z\in\mathcal{Z}}\mathbb{E}_{\mathcal{A}}[|\ell(\mathcal{A(S)};z)-\ell(\mathcal{A(S^{\prime})};z)|]\leq 2{\left(\frac{2C^{2}}{L}\right)}^{\frac{1}{c+1}}{\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}.$
Proof.
@@
For each $t\in\{1,\ldots,T\}$, following the proof sketch in Section D.2, we bound the term $\mathbb{E}_{\mathcal{A}}\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ where $G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}_{t}$, and for each $i\in\mathcal{H}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i,J^{(i)}_{t})}_{t})$ with $J^{(i)}_{t}$ sampled uniformly from $\{1,\ldots,m\}$. We establish the bound by conditioning on whether the differing samples are selected at a given step, where $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ denote the indices of the differing samples. At each iteration, a single sample is selected uniformly at random from each honest worker. Consequently, with probability $(1-\frac{1}{m})$, the differing samples are not selected:
$\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}=\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$
where $\eta^{\mathrm{SGD}}_{G_{\gamma_{t}}}$ denotes the expansive coefficient, which depends on the regularity properties of the loss function (see Lemma B.2). And with probability $\frac{1}{m}$, the differing samples are selected:

### v2

@@ Appendix B REGULARITY ASSUMPTIONS AND EXPANSIVITY RESULTS
To analyze the stability of iterative optimization algorithms, we typically rely on regularity assumptions on the loss function. We begin by stating standard regularity assumptions used throughout the paper, along with classical results that follow from them.
@@
Notably, convex and $L$-smooth functions satisfy an important property known as co-coercivity (see Proposition 5.4 in Bach (2024)). Specifically,
@@
Here, for each $i\in\mathcal{H}$, $g^{(i)}$ denotes the exact gradient $\nabla\ell(\theta)$ in the case of $\mathrm{GD}$, or an unbiased estimate thereof—effectively computed with a sample from the local dataset—in the case of $\mathrm{SGD}$. We recall the expansiveness definition for an update rule from Hardt et al. (2016, Definition 2.3):
@@
A straightforward extension of the result from Hardt et al. (2016, Lemma 3.6) to the distributed setting yields:
@@
Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ and $\gamma>0$. Assume $\forall z\in\mathcal{Z},\ell(\cdot,z):\Theta\to\mathbb{R}$ is L-smooth, Then $G^{\mathcal{A}}_{\gamma}$ is $\eta_{G^{\mathcal{A}}_{\gamma}}$-expansive (Hardt et al., 2016, Definition 2.3), with the following expressions:
@@
Appendix C DEFERED PROOFS FROM SECTION 3.1
We first establish supporting results in Section C.1, then prove Theorem 3.1.
C.1 Expectation bounds for honest gradients’ empirical spectral norm
@@
We derive bounds on the expected spectral norm of the empirical covariance matrix of the honest workers’ gradients under the setting of bounded gradient assumption (i.e. Lipschitz-continuity of the loss function).
Lemma C.1.
@@
Consider the setting described in Section 2 under Byzantine failures. Assume the loss function $C$-Lipschitz. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. We have for every $t\in\{1,\ldots,T-1\}$:
@@
C.2 Proof of Theorem 3.1
We begin with a unified proof sketch that outlines the key ideas to bound stability under Byzantine failures. We then prove Theorem 3.1.
@@
Let denote $\{{\theta_{t}\}}_{t\in\{0,\ldots,T-1\}}$ and ${\{\theta^{\prime}_{t}\}}_{t\in\{0,\ldots,T-1\}}$ the optimization trajectories resulting from two neighboring datasets $\mathcal{S},\mathcal{S}^{\prime}$, for $T$ iterations of $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$ with $F$ a $(f,\kappa)$-robust aggregation rule, and learning rate schedule ${\{\gamma_{t}\}}_{t\in\{0,\ldots,T-1\}}$. Let denote $m^{(i)}_{t}$ and $m^{\prime(i)}_{t}\in\mathbb{R}^{d}$ the local updates provided by worker $i\in\{1,\ldots,n\}$ along the trajectories $t\in\{0,\ldots,T-1\}$. In the context of the uniform algorithmic stability analysis via parameter sensitivity, we track how the parameters diverge along these optimization trajectories. To do so, we decompose the analysis by introducing an intermediate comparison between the robust update $G^{F}_{\gamma_{t}}$ and the averaging over honest workers update $G^{\mathcal{A}}_{\gamma_{t}}$. This enables us to leverage either the regularity of the loss function and the robustness property of the aggregation rule. This approach is primarily motivated by the fact that Byzantine vectors cannot be assumed to exhibit regularities when comparing them.
$\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]$ $=\mathbb{E}_{\mathcal{A}}[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})+G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})+G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}]$
$\leq\underbrace{\mathbb{E}_{\mathcal{A}}\|G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}}_{A}+\underbrace{\gamma_{t}\mathbb{E}_{\mathcal{A}}\left[\|G^{F}_{\gamma_{t}}(\theta_{t})-G^{\mathcal{A}}_{\gamma_{t}}(\theta_{t})\|_{2}+\|G^{F\prime}_{\gamma_{t}}(\theta^{\prime}_{t})-G^{\mathcal{A}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\right]}_{B}$
We then bound the term $A$ with classical stability tools from Hardt et al. (2016), and the term $B$ with the $(f,\kappa)$-robust property, the Jensen inequality and the Lemma C.1:
@@
Theorem 3.1.
Consider the setting described in Section 2 under Byzantine failures. Let $\mathcal{A}\in\{\mathrm{GD},\mathrm{SGD}\}$, with $F$ a $(f,\kappa)$-robust aggregation rule, for $T\in\mathbb{N}^{*}$ iterations. Suppose $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $C$-Lipschitz and $L$-smooth.
@@
    Assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ convex, with constant learning rate $\gamma\leq\frac{2}{L}$, then we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
@@
    Assume $\forall z\in\mathcal{Z}$, $\ell(\cdot;z)$ $\mu$-strongly convex, with constant learning rate $\gamma\leq\frac{1}{L}$, then we have for any neighboring datasets $\mathcal{S},\mathcal{S^{\prime}}$:
@@
Let first reason about $\mathcal{A}=\mathrm{GD}$. For each $t\in\{0,\ldots,T-1\}$, following the proof sketch in C.2, we bound the term $\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ where $G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\nabla\widehat{R}_{\mathcal{H}}(\theta_{t})$. Let $a,b\in\mathcal{H}\times\{1,\ldots,m\}$ denote the indices of the differing samples. For $t\in\{0,\ldots,T-1\}$, we have,
$\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}$ $=\|G^{\mathrm{GD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{GD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}+\frac{\gamma_{t}}{|\mathcal{H}|m}\|\nabla\ell(\theta_{t}^{\prime};z^{\prime(a,b)})-\nabla\ell(\theta_{t}^{\prime};z^{(a,b)})\|_{2}$
@@
where $\eta^{\mathrm{GD}}_{G{\gamma_{t}}}$ denotes the expansive coefficient, which depends on the regularity properties of the loss function (see Lemma B.2). Let define $\sigma^{F}_{\mathrm{GD}}=2C\left(\frac{1}{(n-f)m}+\sqrt{\kappa}\right)$, we then resolve the recursion by incorporating the impact of the Byzantine failures as demonstrated in the previous proof sketch.
@@
For $\mathcal{A}=\mathrm{SGD}$, the same recursion holds for the expected distance $\mathbb{E}_{\mathcal{A}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2}$ for $t\in\{1,\ldots,T\}$, where $G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})=\theta_{t}-\gamma_{t}\frac{1}{|\mathcal{H}|}\sum_{i\in\mathcal{H}}g^{(i)}_{t}$, and for each $i\in\mathcal{H}$, $g^{(i)}_{t}=\nabla\ell(\theta_{t};z^{(i,J^{(i)}_{t})}_{t})$ with $J^{(i)}_{t}$ sampled uniformly from $\{1,\ldots,m\}$. In fact, at each iteration, a single sample is selected uniformly at random from each honest worker. Consequently, with probability $(1-\frac{1}{m})$, the differing samples are not selected,
$\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}\prime}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}=\|G^{\mathrm{SGD}}_{\gamma_{t}}(\theta_{t})-G^{\mathrm{SGD}}_{\gamma_{t}}(\theta^{\prime}_{t})\|_{2}\leq\eta_{G^{\mathrm{SGD}}_{\gamma_{t}}}\|\theta_{t}-\theta^{\prime}_{t}\|_{2},$
and with probability $\frac{1}{m}$, the differing samples are selected,